OpenStudy (anonymous):

If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

@junyang96 can you help me with this?

OpenStudy (anonymous):

If we draw our points like this:  We can tell that since (19, -5) is the midpoint it means that the y-difference between (18, y) and (19,-5) is the same as the y-difference between (19,-5) and (20, -15) which is: $$ -5 - (-15) = -5 + 15 = 10 $$What that means is that the y of (19, -5) is bigger than the y of (20, -15) by 10. So that means that the y of (18, y) is bigger than the y of (19, -5) also by 10: $$ y = -5 + 10 = 5 $$

OpenStudy (anonymous):

My (amazing) drawing was deleted for some reason  \( \Huge ☺ \)

OpenStudy (anonymous):

the answer i came to was that y=5, is that correct? @pitamar

OpenStudy (anonymous):

Ye, that's what I got up here too =)

OpenStudy (anonymous):

Ill tell yo if its correct when the work is graded

OpenStudy (anonymous):

Ok =)